Calculate the molarity when 250 g of C6 H12 O6 is dissolved in 2500 ml of water.


1
Expert's answer
2021-02-03T07:07:10-0500

M(C6H12O6) = 180 g/mol

n(C6H12O6) "= \frac{250}{180} = 1.388 \;mol"

Proportion:

1.388 mol – 2500 mL

x – 1000 mL

"x = \frac{1.388 \times 1000}{2500} = 0.55 \;M"

Answer: 0.55 M